OpenStudy (anonymous):

For the following sequences, find an explicit formula in terms of n for s_n: 1, -(1/2), 1/4, -(1/8), 1/16... 1, 7/9, 10/14, 13/19, 16/24... 1, 3/8, 9/27, 27/64, 81/125...

OpenStudy (anonymous):

With s_n being "s sub n"

OpenStudy (zzr0ck3r):

find the common ratio for the first one that is -1/2 so \(\sum_{n=0}^\infty(-\frac{1}{2})^n\)

OpenStudy (anonymous):

Yeah, but that doesn't work for s_1, because, in that case, it would be -1/2 instead of 1, right?

OpenStudy (zzr0ck3r):

is s_n a series or a sequence?

OpenStudy (zzr0ck3r):

err is it the partial sums of the series ?

OpenStudy (anonymous):

No, it's a sequence. Is that the difference? That in a sequence, it doesn't have to work for the first number in the sequence?

OpenStudy (anonymous):

By the way, these make absolutely no sense to me, so if I'm missing something remedial, call me out on it.

OpenStudy (zzr0ck3r):

\(s_n = (-\frac{1}{2})^{n-1}=\{1,-\frac{1}{2},\frac{1}{4},-\frac{1}{8}......\}\)

OpenStudy (zzr0ck3r):

normally we reserve \(s_n=\sum_{i=1}^n a_n\) where \(a_n \) is a sequence.

OpenStudy (anonymous):

Ah. My calculus teacher has s_n on our worksheet.

OpenStudy (zzr0ck3r):

and you have not started to study series yet?

OpenStudy (anonymous):

I don't believe so. There's nothing of series on our worksheet. Then again, our teacher is all over the place at any given time during lecture, so who knows.

OpenStudy (zzr0ck3r):

I'm just trying to make sure I'm telling you the right thing...

OpenStudy (zzr0ck3r):

you would know

OpenStudy (anonymous):

I believe that series are tomorrow's lecture.

OpenStudy (anonymous):

Yes, Series are on Wednesday. Just checked the syllabus.

OpenStudy (zzr0ck3r):

first we start with sequences \(a_n=a_1,a_2, a_3.....\) then we start adding them up, and we say \(s_n=a_1+a_2+a_3+....+a_n\)

OpenStudy (zzr0ck3r):

ok then forget the first thing I said and go with the second:)

OpenStudy (anonymous):

Alright, well, I knew it had to do something with -1/2 from the start, but these other ones I'm completely and utterly lost on.

OpenStudy (anonymous):

That's my issue: I can't seem to think of these kind of equations out of the blue.

OpenStudy (zzr0ck3r):

hmm the 1 is throwing me, but im not the best at these...i just try and construct it.

OpenStudy (zzr0ck3r):

if the 1 was not there then we could use \(s_n = (\frac{7+3(n-1)}{9+5(n-1)})\)

OpenStudy (anonymous):

I can see a pattern, like, "Ok, the numerator is increasing by 3 each time, and the denominator is increasing by 5 each time" but after that I have nothing.

OpenStudy (zzr0ck3r):

thats exactly how I came up with that^^^

OpenStudy (zzr0ck3r):

hmm I got to crash because its 1am and I got a final in 8 hours:( but gl.

OpenStudy (anonymous):

Alright, good luck and thanks!

OpenStudy (kc_kennylau):

Try to describe the pattern using words first, then I'd teach you how to convert words into mathematical notation

OpenStudy (anonymous):

Alright, here's my go at the second sequence: The numerator is increasing by 3 for each iteration, and the denominator is increasing by 5 for each iteration.

OpenStudy (kc_kennylau):

Yep, try to do the numerator and the denominator separately.

OpenStudy (anonymous):

Alright, give me a sec.

OpenStudy (kc_kennylau):

Okay, take your time!

OpenStudy (anonymous):

\[(4+3(n-1)) / (4+5(n-1))\]

OpenStudy (anonymous):

I'm not very good with the equation editor.

OpenStudy (kc_kennylau):

Then simplify it.

OpenStudy (kc_kennylau):

You don't need to use the equation editor, just typing it out would do the job

OpenStudy (anonymous):

Alright, but as a neat freak, I like it when I can make the equations look nice for others to see. I'll simplify it here in a sec.

OpenStudy (kc_kennylau):

(4+3(n−1))/(4+5(n−1)) is as nice as \(\dfrac{4+3(n-1)}{4+5(n-1)}\) :)

OpenStudy (anonymous):

I ended up with \[\frac{ 1+3n }{ 5n-1 }\]

OpenStudy (kc_kennylau):

Perfect!

OpenStudy (anonymous):

YES. I GOT IT TO WORK. (The editor)

OpenStudy (kc_kennylau):

P.S. I don't use the editor

OpenStudy (anonymous):

Now, about this third one, I got the numerator, but the denominator is throwing me off.

OpenStudy (kc_kennylau):

Hint: cubes

OpenStudy (anonymous):

Alright. Gimmi a minute to look at it again.

OpenStudy (kc_kennylau):

Need the answer?

OpenStudy (anonymous):

Nonono, like, one more minute!

OpenStudy (kc_kennylau):

Okay, take your time!

OpenStudy (anonymous):

\[\frac{ 3^{n-1} }{ n ^{3} }\]

OpenStudy (anonymous):

?

OpenStudy (kc_kennylau):

Perfect!

OpenStudy (kc_kennylau):

Take note of the result generated by the equation editor every time you use it

OpenStudy (kc_kennylau):

Soon you'll be able to type equations without the editor!

OpenStudy (anonymous):

Yeah, I'm sure if I use it more, it'll be nicer to not have to use it, but for now, I'll take my crutch and run with it.

OpenStudy (anonymous):

Err, hobble, maybe.

OpenStudy (anonymous):

Thanks for your help, and teaching me what took my teacher a few days, in roughly a half hour or so!

OpenStudy (kc_kennylau):

Wow, I'm flattered. No problem!